A mass M is split into two parts, m and M-m, which are then separated by a certain distance. What ratio m/M maximizes the magnitude of the gravitational force between the pairs?

Grade:6

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Nidhi

F=Gm(M-m)/r2

dF/dm = G/r2 (M-2m)=0

        m=M/2

m/M =1/2 

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin